# Reduced Order Hysteretic Magnetization Model for Composite Superconductors

### v1

@@ 6.1 Parameter identification
The reference solution is the major loop represented in Fig. 15. We choose $N=6$ and apply the two-step procedure described in Section 5.1 with $N$ equally spaced values $\mu_{0}\bar{\kappa}_{k}$ from 00 T to $0.75$ T. The obtained parameters are given in Table 2.
$k$ $\alpha_{k}$ $\mu_{0}\bar{\kappa}_{k}$
@@
$1$ $48.24$ 00
$2$ $18.07$ $150$
$3$ $16.99$ $300$
$4$ $9.31$ $450$
$5$ $5.16$ $600$
$6$ $2.22$ $750$
Table 2: Parameters of a chain of S cells with $N=6$, identified on the reference magnetization loop shown in Fig. 15.
It is interesting to notice that the first S cell, which is anhysteretic, almost contributes to half of the magnetic flux density. This is due to the large fraction of normal conductor in the strand, the matrix, which does not behave as a hysteretic material.
@@ 6.2 Unidirectional excitation
Using the parameters obtained above, we compare the predictions of the hysteresis model in terms of magnetization and loss for two unidirectional excitations along $\hat{\boldsymbol{e}}_{y}$. The first one is a harmonic field:
@@
with $\mu_{0}h_{\text{max}}=2$ T and $f=0.01$ Hz. The second one is a biharmonic field, so as to produce minor magnetization loops:
@@
Results are given in Fig. 16. With $N=6$ cells, the hysteresis model is capable of reproducing faithfully both major and minor magnetization loops. The loss is accurately reproduced in both cases, with a relative error on the total loss below $1\%$.
The influence of the number of cells on the relative error on total loss is shown in Fig. 17. For the considered excitations, increasing the number of cells helps reducing the error for low $N$ values ($\leq 3$), but the error then stabilizes to a non-zero value. One source of error is related to the scaling function $f_{\kappa}(\boldsymbol{b})$ for irreversibility parameters, which is only approximate. Indeed, in reality, the magnetic flux density $\boldsymbol{b}$ is not uniform in the filament, but rather varies due to the screening currents and geometrical effects. Also, as mentioned above, the reference solution is not truly rate-independent, due to the finite $n$ value in the power law.
Figure 17: Evolution of the relative error $(Q_{\text{hyst}}-Q_{\text{ref}})/Q_{\text{ref}}$ between total losses predicted by the hysteresis model ($Q_{\text{hyst}}$) and the reference model ($Q_{\text{ref}}$), for different sources, as a function of the number of cells $N$, with equally spaced values $\mu_{0}\bar{\kappa}_{k}$, from 00 T to $0.75$ T. In the case $N=1$, $\mu_{0}\bar{\kappa}_{1}=0.168$ T and $\alpha_{1}=1$.
The major advantage of the method is its computational efficiency. Once the model parameters are identified, solving the hysteresis model is tremendously faster than solving a finite element model (typically a few milliseconds compared to a few tens of minutes). This efficiency is crucial in view of homogenizing full-scale superconducting systems.
In practice, the number of cells and the distribution of $\bar{\kappa}_{k}$ values must be chosen depending on the actual excitations to be considered in the end application. A chain of S cells produces no loss for field variations smaller than the smallest non-zero irreversibility parameter $\kappa_{k}(\boldsymbol{b})$, as illustrated in Fig. 18. To reproduce accurately power loss at low fields, or for low field ripples, the chain of S cells must therefore contain cells with sufficiently small irreversibility parameters. This will be further illustrated in the case of a chain of CS cells in Section 7.
Figure 18: Illustration of a chain of S cells with $N=3$. The centers of the spheres are at positions $\boldsymbol{h}_{\text{rev},k}$, established by the history of $\boldsymbol{h}$. Variations of the magnetic field $\boldsymbol{h}$ that stay inside the smallest sphere do not produce any loss.
@@ 6.3 Bidirectional excitation
The parameters were identified based on the results of a unidirectional excitation, but all the hysteresis model equations are vectorial, and the model is therefore directly applicable to general excitations. We show in Fig. 19 the results for a bidirectional rotating transverse field excitation defined by:
@@
Figure 19: Comparison of hysteresis model results with a reference solution. Rotating transverse magnetic field $\boldsymbol{h}_{\text{app}}^{(3)}(t)$. Chain of S cells with $N=6$ cells and parameters of Table 2. Solutions at selected time instants are indicated by the colored circles (reference solution) and crosses (hysteresis model). (Up) Internal magnetic field and magnetization of the filament (same legend as in the bottom sub-figure). Values are in tesla (T). (Down) Dissipated power per unit length as a function of time.
The hysteresis model reproduces well the magnetization and power loss. The angle between the inner field $\boldsymbol{h}_{\text{in}}$ and the magnetization vector $\boldsymbol{m}$ is faithfully described, as shown by the solutions at selected time instants in Fig. 19. The agreement is not perfect, but very satisfying provided that the model parameters were fully identified using results from a unidirectional situation, with absolutely no information about the strand response in more general cases. We let the questions of improving the agreement and tackling anisotropic systems for further works.
@@ 7 Results with a chain of CS cells
As a generalization of the previous section, we consider the same 54-filament strand subject to a transverse field, but now with higher rates of field changes, so that both magnetization and loss exhibit frequency dependence. The reference solutions are obtained as described in Section 2 with the CATI method in order to account for coupling current effects.
We start the analysis by choosing the hysteresis model parameters based on reference solutions. We then illustrate the results for wide ranges of field amplitudes and frequencies.
@@ 7.1 Parameter identification
We choose the chain structure represented in Fig. 14. Compared to the chain of S cells of the previous section, two cells now have a zero irreversibility parameter, $\kappa_{1}=0$ A/m and $\kappa_{2}=0$ A/m. They allow to reproduce losses at low fields, for which filament loss is small with respect to coupling and eddy loss. These two cells allow to reproduce the transitions between the three regimes illustrated in Fig. 20 at low fields, associated with different magnetization slopes, as shown in Fig. 21. The curves are obtained with the parameter values of Table 3, which are discussed below.
Figure 20: Field distribution at low fields ($\mu_{0}h=10$ mT here) and different frequencies. The lines are parallel to the local magnetic field, the larger and darker they are, the higher the field amplitude. (a) Uncoupled filament and limited magnetization. (b) Increased magnetization due to coupled filaments. (c) Close to perfect diamagnetism due to eddy currents.
Figure 21: Strand magnetization for $\mu_{0}h_{\text{max}}=10$ mT at different frequencies, from virgin state at time $t=0$ to time $t=0.25/f$ (for better clarity). Dashed lines are reference solutions and solid lines are from the hysteresis model.
At low frequencies, the dominant contribution to magnetization and loss comes from superconducting filament hysteresis. We can reuse the material parameters found in the previous section for the chain of S cells as was given in Table 2. In order to reproduce losses at fields lower than $0.15$ T, we also introduce $8$ new cells with lower irreversibility parameters, logarithmically spaced from $1$ mT to $80$ mT. The associated weights can be identified as proposed in Section 5.1, or based on analytical solutions at low fields.
@@
Finally, we fix the scaling $f_{\chi}(\boldsymbol{b})$ and the parameters $\bar{\chi}_{k}$. We found that a decent fit is obtained with
$f_{\chi}(\boldsymbol{b})=\frac{1-\|\boldsymbol{b}\|/b_{0}}{1+\|\boldsymbol{b}\|/b_{1}},$ (45)
with $b_{0}=15$ T and $b_{1}=4$ T. The parameters $\bar{\chi}_{k}$ are then tuned manually to obtain the values in Table 3.
$k$ $\alpha_{k}$ $\mu_{0}\bar{\kappa}_{k}$ $\tau_{\text{e},k}$ $\tau_{\text{c},k}$ $\mu_{0}\bar{\chi}_{k}$
@@ 7.2 Results and numerical performance
Using the model parameters given in Table 3, we compare the predictions of the hysteresis model with the reference solutions. The power loss per cycle as a function of frequency and selected magnetization loops are given in Figs. 22 and 23. The power loss per cycle is also presented as a map in Fig. 24, by interpolating the results of $400$ simulations, spanning over $25$ frequencies and $16$ field amplitudes, logarithmically spaced.
Figure 22: Power loss per cycle for three different field amplitudes as a function of frequency. Solid curves are results from the hysteresis model with parameters of Table 3 and dashed curves are reference solutions. The legend is the same for the three subfigures.
Figure 23: Strand magnetization for $\mu_{0}h_{\text{max}}=1$ T at different frequencies. Solid curves are results from the hysteresis model with parameters of Table 3 and dashed curves are reference solutions.
Figure 24: Total loss per cycle and per unit length as a function of the frequency and amplitude of the applied field. Dashed curves are contour lines of equal loss. Solid curves delimit areas of distinct dominant loss contributions (from left to right: hysteresis in uncoupled filaments, coupling loss, hysteresis in coupled filaments, and eddy losses). (Left) Reference model, adapted from [31]. (Right) Prediction of the hysteresis model with the parameters given in Table 3.
The overall agreement on the total loss and magnetization is very good. The hysteresis model correctly reproduces the different regimes of distinct dominant loss contributions and describes the magnetization faithfully. The difference with the reference solution is higher at high frequency and high fields, as can be seen on the contour lines in the top right corners in Fig. 24. This can be improved with a careful choice of model parameters, but as discussed in Section 3.3, the matching will never be perfect in the limit $f\to\infty$ as the CS cells do not reproduce a $1/\sqrt{f}$ decrease of the loss per cycle, but rather a $1/f$ decrease. For not too high frequencies however, the importance of this effect is limited.
The major advantage of the hysteresis model compared to the detailed simulations is the very small associated computational work. Even with $N=15$ cells, the whole loss map on the right of Fig. 24, which consists of $400$ hysteresis model simulations, can be performed in a few seconds, whereas the reference model requires $26$ days of computation time in total to generate the loss map on the left of Fig. 24.
@@
In superconducting magnets, many strands are combined together into cables, which themselves are wound around the magnet aperture. The hysteresis model can be used to homogenize the large number of strands into an equivalent homogeneous material. In such a case, which is let for further work, the reference detailed solution should be that of a close-packed set of strands, which can typically be modelled via appropriate periodic boundary conditions [14], rather than a single strand in air, as was done here for the sake of illustration. This may result in different model parameters, but the general approach is unchanged.
@@ 8 Conclusion
In this work, we introduced the Reduced Order Hysteretic Magnetization (ROHM) model to represent the magnetization and instantaneous loss in composite superconductors subject to transient external fields. This model is designed by adapting to superconducting systems the state-of-the-art energy-based model developed in the context of ferromagnetic hysteresis [27]. We focused specifically on the case of a multifilamentary superconducting strand and proposed two models: a rate-independent model, describing hysteresis in superconducting filaments, and a rate-dependent model, accounting for filament coupling, coupling current, and eddy current effects.
We proposed a parameter identification approach for both models and described their inclusion in a finite element framework. Finally, we demonstrated that these models help to strongly reduce the computational cost compared to conventional simulations, while offering a very good accuracy in a wide range of field amplitudes and rates of field change.
This work constitutes a first step towards the homogenization of superconducting magnets, in which the large number of turns makes detailed simulations unrealistically expensive. Replacing the coil windings by homogenized material properties such as the proposed ROHM model allows to describe magnetization and loss efficiently. Further work is necessary to include non-zero transport current, which was not covered in this paper.
Extension of the method to anisotropic systems such as high-temperature superconducting tapes or stacks of tapes in view of their homogenization can also be considered as further work.
@@ Appendix A Magnetization in 2D Problems
The magnetic dipole moment $\boldsymbol{M}$ of a current density distribution $\boldsymbol{j}$ in a given volume $V$ is evaluated as
@@
with $\boldsymbol{x}$ the position vector. Consider a rectangular current tube of cross section $\text{d}S$, width $w$, and length $l$, carrying a current $\text{d}I=\boldsymbol{j}\cdot\text{d}\boldsymbol{S}$, as illustrated in Fig. 26. Accounting for the four sides of the rectangle, the contribution of this current loop to the total magnetic dipole moment is $wl\,\text{d}I$. Long and short sides of the rectangle contribute equally to the total magnetic dipole moment.
Figure 26: Magnetic dipole moment of a current tube in a cylinder.
When the geometry is sufficiently long for end effects to be neglected, the problem is usually solved in 2D and per unit length, on a cross section of the conducting cylinder. In such a case, however, one must still account for the fact that all current loops close at the end of the cylinder, i.e., at infinity, for computing the magnetization per unit length. These closing currents are not part of the 2D solution, and are then missing from any magnetization calculation performed solely on the 2D cross section. Because the closing currents contribute to exactly half of the magnetic dipole moment, one can account for them by introducing a factor $2$, as is done in Eq. (1).

### v2

@@ 6.1 Parameter identification
The reference solution is the major loop represented in Fig. 14. We choose $N=6$ and apply the two-step procedure described in Section 5.1 with $N$ equally spaced values $\mu_{0}\bar{\kappa}_{k}$ from 00 mT to $750$ mT. The obtained parameters are given in Fig. 17.
$k$ $\mu_{0}\bar{\kappa}_{k}$ $\alpha_{k}$
@@
$1$ 00 $48.24$
$2$ $150$ $18.07$
$3$ $300$ $16.99$
$4$ $450$ $9.31$
$5$ $600$ $5.16$
$6$ $750$ $2.23$
Figure 17: (Left) Parameters of a chain of S cells with $N=6$, identified on the reference magnetization loop shown in Fig. 14. (Right) Scaling function $f_{\kappa}(b)$.
It is interesting to notice that the first S cell, which is anhysteretic ($\kappa_{1}=0$ A/m), almost contributes to half of the magnetic flux density. This is due to the large fraction of normal conductor in the strand, the copper matrix (copper fraction in the cross section is $56\%$), which does not behave as a hysteretic material.
(a) Major loop $h_{\text{app}}^{(1)}(t)$.
(b) Minor loops $h_{\text{app}}^{(2)}(t)$.
Figure 18: Comparison of the ROHM model results with the reference solution. Unidirectional transverse magnetic field. Chain of S cells with $N=6$ cells, parameters of Fig. 17. (a) Harmonic excitation $h_{\text{app}}^{(1)}(t)$. (b) Biharmonic excitation $h_{\text{app}}^{(2)}(t)$. (Up) Internal magnetic field, the input of the ROHM model. (Middle) Magnetization loop. The curves of the ROHM model are colored as a function of the instantaneous power, see the bottom plots for the legend. (Down) Dissipated power per unit length as a function of time, equal to $a\,p_{\text{irr}}(t)$, with $a=\pi d^{2}/4$ the surface area of the strand cross section.
@@ 6.2 Unidirectional excitation
Using the parameters obtained above, we compare the predictions of the ROHM model in terms of magnetization and loss for two types of unidirectional excitations along $\hat{\boldsymbol{e}}_{y}$. The first one is a harmonic field:
@@
with $\mu_{0}h_{\text{max}}=2$ T and $f=0.01$ Hz. The second one is a biharmonic field, resulting in a response with minor magnetization loops:
@@
Results are given in Fig. 18. With $N=6$ cells, the ROHM model reproduces very well both major and minor magnetization loops. The instantaneous loss is remarkably well reproduced in both cases, with a relative error on the total loss below $1\%$.
The influence of the number of cells on the relative error on the total loss is shown in Fig. 19. Increasing the number of cells helps reducing the error for low $N$ values ($\leq 3$). The error then stabilizes (for $N\geq 6$) to a small, non-zero value.
Figure 19: Evolution of the relative error $(Q_{\text{hyst}}-Q_{\text{ref}})/Q_{\text{ref}}$ between total losses predicted by the ROHM model ($Q_{\text{hyst}}$) and the reference model ($Q_{\text{ref}}$), for different excitations, as a function of the number of cells $N$, with equally spaced values $\mu_{0}\bar{\kappa}_{k}$, from 00 T to $0.75$ T. In the case $N=1$, $\mu_{0}\bar{\kappa}_{1}=0.168$ T and $\alpha_{1}=1$.
The major advantage of the method is its computational efficiency. Once the model parameters are identified, solving the ROHM model is tremendously faster than solving a detailed FE model. Typically a few milliseconds only is required for the ROHM model compared to a few tens of minutes for the 54-filament reference FE simulation. This efficiency is crucial in view of simulating full-scale superconducting systems.
In practice, the number of cells and the distribution of $\bar{\kappa}_{k}$ values must be chosen depending on the actual excitations to be considered in practice. As already mentioned, a chain of S cells produces no loss for field variations smaller than the smallest non-zero irreversibility parameter $\kappa_{k}(\boldsymbol{b})$, as illustrated in Fig. 15. To reproduce accurately power loss at low fields, or for low field ripples, the chain of S cells must therefore contain cells with sufficiently small irreversibility parameters. The same comment holds for a chain of CS cells, and will be further illustrated in Section 7.
@@ 6.3 Bidirectional excitation
The parameters in Section 6.1 are defined based on the results of a unidirectional excitation, but all the ROHM model equations are vectorial, and the model is therefore directly applicable to general excitations. We show in Fig. 20 the results for a bidirectional rotating transverse field excitation defined by:
@@
Figure 20: Comparison of ROHM model results with the reference solution. Rotating transverse magnetic field $\boldsymbol{h}_{\text{app}}^{(3)}(t)$. Chain of S cells with $N=6$ cells and parameters of Fig. 17. Solutions at selected time instants are indicated by the colored circles (reference solution) and crosses (ROHM model). (Up) Internal magnetic field and magnetization of the filament. The ROHM model curve is colored as a function of the instantaneous power loss. Same legend as in the bottom sub-figure. Values are in tesla (T). (Down) Dissipated power per unit length as a function of time.
The ROHM model reproduces well the magnetization and power loss. The angle between the inner field $\boldsymbol{h}_{\text{in}}$ and the magnetization vector $\boldsymbol{m}$ is faithfully described, as shown by the solutions at selected time instants in Fig. 20. The agreement is not perfect, but very satisfying provided that the model parameters are fully identified using results from a unidirectional situation, with absolutely no information about the strand response under bidirectional excitations.
@@ 7 Results with a chain of CS cells
As a generalization of the previous section, we consider the same 54-filament strand subject to a transverse field, but now with a conducting matrix and higher rates of field changes, so that both magnetization and loss exhibit frequency dependence. The reference solutions are obtained as described in Section 2 with the CATI method in order to account for coupling current effects due to the twist of the filaments.
We start the analysis by choosing the ROHM model parameters based on reference solutions. We then illustrate the results for wide ranges of field amplitudes and frequencies.
@@ 7.1 Parameter identification
We choose the chain structure represented in Fig. 21. Compared to the chain of S cells of the previous section, two cells now have a zero irreversibility parameter, $\kappa_{1}=0$ A/m and $\kappa_{2}=0$ A/m. These two cells allow to reproduce the transitions between the three regimes illustrated in Fig. 22 at low fields, associated with different magnetization slopes, as shown in Fig. 23. The curves are obtained with the parameter values of Table 1, which are discussed below.
Figure 21: Chain of CS cells to model a multifilamentary strand rate-dependent response.
Figure 22: Field distribution at low fields ($\mu_{0}h_{\text{max}}=10$ mT) and different frequencies. The lines are parallel to the local magnetic field, the larger and darker they are, the higher the field amplitude. (a) Uncoupled filaments and limited magnetization. (b) Increased magnetization due to coupled filaments. (c) Almost perfect diamagnetism due to eddy currents.
Figure 23: Strand magnetization for $\mu_{0}h_{\text{max}}=10$ mT at different frequencies, from virgin state at time $t=0$ to time $t=0.25/f$ (for better clarity). Dashed lines are reference solutions and solid lines are from the ROHM model.
At low frequencies, the dominant contribution to magnetization and loss comes from superconducting filament hysteresis. We can reuse the material parameters found in the previous section for the chain of S cells as is given in Fig. 17. In order to reproduce losses at fields lower than $150$ mT, we also introduce $8$ new cells with lower irreversibility parameters, logarithmically spaced from $1$ mT to $80$ mT. The associated weights can be identified automatically as proposed in Section 5.1, or based on analytical solutions at low fields. These additional cells are optional and only needed to reproduce accurate results on wide field amplitude ranges.
@@
Finally, we fix the scaling $f_{\chi}(\boldsymbol{b})$ and the parameters $\bar{\chi}_{k}$. A decent fit is obtained with the curve in Fig. 24. The parameters $\bar{\chi}_{k}$ are then tuned manually to obtain the values in Table 1.
Figure 24: Field-dependent scaling functions for the irreversibility parameters $\kappa_{k}(b)$ and $\chi_{k}(b)$ for the chain of CS cells.
$k$ $\mu_{0}\bar{\kappa}_{k}$ $\alpha_{k}$ $\tau_{\text{e},k}$ $\tau_{\text{c},k}$ $\mu_{0}\bar{\chi}_{k}$
@@ 7.2 Results and numerical performance
Using the model parameters given in Table 1, we compare the predictions of the ROHM model with the reference solutions. The loss per cycle as a function of frequency and magnetization loops are given in Figs. 25 and 26.
Figure 25: Energy loss per cycle for three different field amplitudes as a function of frequency. Solid curves are results from the ROHM model with parameters of Table 1 and dashed curves are results from the reference solutions. The legend is the same for the three subfigures.
Figure 26: Strand magnetization for $\mu_{0}h_{\text{max}}=1$ T at different frequencies. Solid curves are results from the ROHM model with parameters of Table 1 and dashed curves are results from the reference solutions.
The overall agreement on the total loss and magnetization is very good. The ROHM model correctly reproduces the different regimes of distinct dominant loss contributions and describes the magnetization faithfully. The difference with the reference solution is larger at high frequency and high fields, as can be seen in Fig. 25. This can be improved with an adapted choice of model parameters, but as discussed in Section 3.3, the matching will never be perfect in the limit $f\to\infty$ as the CS cells do not reproduce a $1/\sqrt{f}$ decrease of the loss per cycle, but rather a $1/f$ decrease. For not too high frequencies however (skin depth not too small compared to strand diameter), the importance of this effect is limited.
As in the rate-independent case, the major advantage of the ROHM model compared to the detailed simulations is the very small associated computational work. Even with $N=15$ cells, each ROHM model simulation is conducted in a few milliseconds, whereas the computational time for a reference solution with the CATI method is of the order of one hour (with 28 0002800028\,00028 000 degrees of freedom). The ROHM model enables the simulation of large-scale superconducting systems involving hundreds or thousands of strands, which would be completely unattainable with detailed strand models.
@@
In superconducting magnets, many strands are combined together into cables, which themselves are wound around the magnet aperture. The ROHM model can be used to homogenize the large number of strands into an equivalent homogeneous material. In such a case, the reference detailed solution should be that of a close-packed set of strands, which can typically be modelled via appropriate periodic boundary conditions [14], rather than a single strand in air, as is done here for the sake of illustration. This may result in different model parameters, but the general approach remains the same.
@@ 8 Conclusion
In this work, we introduced the Reduced Order Hysteretic Magnetization (ROHM) model to represent the magnetization and instantaneous power loss in composite superconductors, subject to transient transverse external fields. This model is designed by adapting the state-of-the-art energy-based model developed in the context of ferromagnetic hysteresis [27, 28].
We focused specifically on the case of a multifilamentary superconducting strand and proposed two models: a rate-independent model, describing hysteresis in superconducting filaments, and a rate-dependent model, accounting for filament coupling, coupling current, and eddy current effects. We proposed a parameter identification approach for both models and described their inclusion in a finite element framework. Finally, we demonstrated that these models have the potential to reduce the computational cost compared to conventional simulations by several orders of magnitude. The model accuracy can be tailored for specific applications. A very good accuracy (typically with a few percent error on the power loss) in a wide range of field amplitudes and rates of field change can be obtained if required.
This work constitutes a first step towards the homogenization of superconducting magnets, in which the large number of turns makes detailed simulations unrealistically expensive. Replacing the coil windings by homogenized material properties such as the proposed ROHM model allows to describe magnetization and loss efficiently. Further work is necessary to include non-zero transport current, which is not covered in this paper.
Extension of the method to anisotropic systems such as high-temperature superconducting (HTS) tapes, stacks of HTS tapes, or HTS bulks in view of their homogenization can also be considered as further work.
@@ Appendix A Magnetization in 2D Problems
The magnetic dipole moment $\boldsymbol{M}$ (A m2) of a current density distribution $\boldsymbol{j}$ in a given volume $V$ is evaluated as
@@
with $\boldsymbol{x}$ the position vector. Consider a rectangular current tube of cross section $\text{d}S$, width $w$, and length $l$, carrying a current $\text{d}I=\boldsymbol{j}\cdot\text{d}\boldsymbol{S}$, as illustrated in Fig. 28. Accounting for the four sides of the rectangle, the contribution of this current loop to the total magnetic dipole moment is $wl\,\text{d}I$. The two long sides of length $l$ contribute to half of it: $wl\,\text{d}I/2$. The two short sides of length $w$, closing the current loop, contribute to exactly the other half of it: $wl\,\text{d}I/2$.
Figure 28: Current carrying rectangular tube in a cylinder.
When the geometry is sufficiently long for end effects to be neglected, the problem is usually solved in 2D, per unit length, on a cross section of the conducting cylinder. To compute the magnetic dipole moment in such a case, however, one must still account for the fact that all current loops must close at the end of the cylinder. These closing currents are not part of the 2D solution, and are then missing from any calculation performed solely on the 2D cross section. Because the closing currents conveniently contribute equally to the magnetic dipole moment compared to the other currents, one can account for them by introducing a factor $2$, as in Eq. (1).